A first year undergraduate statistics course consisting of 1000 students (in different sections) has a mean average of 70.2% with a standard deviation of 3.2%. Only students with a mark of 75% or greater are accepted into the Honours program and continue into second year. How many students will have a mark high enough to move on?
Answer by Boreal(15235)

z=(x-mean)/sd
=(75-70.2)/3.2=4.8/3.2=+1.5
probability z>1.5 is 0.0668
That would be 66.8 students, 67 if rounded up, 66 if rounded down.